OpenStudy (anonymous):

Can someone walk me through: 6^(2x)=3 square root of 36

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[6^{2x}=\sqrt[3]{36}\]

OpenStudy (danjs):

ln(x^b) = b* ln(x)

OpenStudy (anonymous):

is this is the Q?

OpenStudy (anonymous):

anyway take the (ln) for both side\[\ln(6^{2x})=\ln(\sqrt[3]{36})\]

OpenStudy (anonymous):

\[2x \ln(6)=\ln(\sqrt[3]{36)}\] \[x=\frac{ \ln(\sqrt[3]{36} )}{ 2\ln(6)}\]

OpenStudy (anonymous):

\[5(2^{2x})+17(2^{x})=-6\]